# A Comprehensive Survey of Dataset Distillation

### v2

@@ Abstract
Deep learning technology has developed unprecedentedly in the last decade and has become the primary choice in many application domains. This progress is mainly attributed to a systematic collaboration in which rapidly growing computing resources encourage advanced algorithms to deal with massive data. However, it has gradually become challenging to handle the unlimited growth of data with limited computing power. To this end, diverse approaches are proposed to improve data processing efficiency. Dataset distillation, a dataset reduction method, addresses this problem by synthesizing a small typical dataset from substantial data and has attracted much attention from the deep learning community. Existing dataset distillation methods can be taxonomized into meta-learning and data matching frameworks according to whether they explicitly mimic the performance of target data. Although dataset distillation has shown surprising performance in compressing datasets, there are still several limitations such as distilling high-resolution data. This paper provides a holistic understanding of dataset distillation from multiple aspects, including distillation frameworks and algorithms, factorized dataset distillation, performance comparison, and applications. Finally, we discuss challenges and promising directions to further promote future studies on dataset distillation.
@@ 1 Introduction
During the past few decades, deep learning has achieved remarkable success in a wide range of applications, including computer vision [57, 41, 19, 42, 20], neural language processing [55, 112, 58], autonomous vehicles [39, 1, 31], and protein structure prediction [100, 51]. An important reason for this success is due to powerful computing resources, which allow deep neural networks to directly handle giant datasets and bypass complicated manual feature extraction, which causes potential loss of data information [109, 110]. For example, the powerful large language model GPT-3 contains $175$ billion parameters and is trained on $45$ terabytes of text data with thousands of graphics processing units (GPUs) [8]. However, massive data are generated every day [97], which pose a significant threat to training efficiency and data storage, and deep learning might reach a bottleneck due to the mismatch between the volume of data and computing resources [104]. Recently, many methods have been proposed to improve training efficiency in deep learning from several perspectives as follows:
  • Quantization: This approach sacrifices data at the byte level during the training process for acceleration [40, 82, 30, 36].
  • Model pruning: This approach removes trainable parameters that have little influence on the final performance [107, 76, 2, 75].
  • Optimization: This approach designs the training algorithms for fast convergence [125] or less memory cost [4, 52].
  • Dataset reduction: this approach generates few representative data to constitute a new training set. Depending on whether the generated data are natural or synthetic, dataset reduction can be classified into coreset selection [83, 22] and dataset distillation [114].
@@
The above narrative shows that efficient deep learning is an extensive task and is beyond the scope of this paper; readers can refer to [108, 81] for more details. In this survey, we focus on dataset distillation to improve training efficiency by synthesizing training data. To better appreciate dataset distillation, we briefly introduce other cognate methods in dataset reduction. For a support vector machine (SVM), its hyperplane is solely determined by “support vectors”, and removing all other points in the training dataset does not influence on the convergence result [23]. Therefore, the selection of “support vectors” is a favorable method to reduce the training burden for SVMs [84]. When the scenario expands to deep learning, the problem of selecting “support vectors” generalizes to the well-known coreset selection; this is the algorithm selects a few prototype examples from the original training dataset as the coreset, and then the model is solely trained on the small coreset to save training costs while avoiding large performance drops. However, elements in the coreset are unmodified and constrained by the original data, which considerably restricts the coreset’s expressiveness, especially when the coreset budget is limited. Recently, a novel approach, dataset distillation (DD)  [^1] Dataset distillation is also referred to as dataset condensation in some studies., has attracted growing attention from the deep learning community and has been leveraged in many practical applications [79, 70, 16, 17]. Different from coreset selection, dataset distillation removes the restriction of uneditable elements and carefully modifies a small number of examples to preserve more information, as shown in Figure 4 for synthetic examples. By distilling the knowledge of the large original dataset into a small synthetic set, models trained on the distilled dataset can acquire a comparable generalization performance. A general illustration for dataset distillation is presented in Figure 1.
Due to the property of extremely high dimensions in the deep learning regime, the data information is hardly disentangled to specific concepts, and thus distilling numerous high-dimensional data into a few points is not a trivial task. Based on the objectives applied to mimic target data, dataset distillation methods can be grouped into meta-learning frameworks and data matching frameworks, and these techniques in each framework can be further classified in a much more detailed manner. In the meta-learning framework, the distilled data are considered hyperparameters and optimized in a nested loop fashion according to the distilled-data-trained model’s risk w.r.t. the target data [77, 114]. The data matching framework updates distilled data by imitating the influence of target data on model training from parameter or feature space [129, 13, 128]. Figure 3 presents these different categories of DD algorithms in a tree diagram.
Figure 2: The schematic structure of dataset distillation and the relationship between the adjacent sections. The body of this survey mainly contains the fundamentals of dataset distillation, taxonomy of distillation schemes, types of factorized $D$, distillation algorithms, performance comparison, applications, challenges, and future directions. Note that ‘Section’ is abbreviated as ‘Sec.’ in this figure.
@@
Apart from directly considering the synthetic examples as optimization objectives, in some studies, a proxy model is designed, which consists of latent codes and decoders to generate highly informative examples and to resort to learning the latent codes and decoders. For example, Such et al. [105] employed a network to generate highly informative data from noise and optimized the network with the meta-learning framework. Zhao and Bilen [127] optimized the vectors and put them into the generator of a well-trained generative adversarial network (GAN) to produce the synthetic examples. Moreover, Kim et al. [54], Deng and Russakovsky [26], Liu et al. [71], Lee et al. [62] learned a couple of latent codes and decoders, and then synthetic data were generated according to the different combinations of latent codes and decodes. With this factorization of synthetic data, the compression ratio of DD can be further decreased, and the performance can also be improved due to the intraclass information extracted by latent codes.
In this paper, we present a comprehensive survey on dataset distillation, and the main objectives of this survey are as follows: (1) present a clear and systematic overview of dataset distillation; (2) review the recent progress and state-of-the-art algorithms and discuss various applications in dataset distillation; (3) give a comprehensive performance comparison w.r.t. different dataset distillation algorithms; and (4) provide a detailed discussion on the limitation and promising directions of dataset distillation to benefit future studies. Recently, Larasati et al. [59] published a short review on dataset distillation. However, these authors only provided a general overview of DD and focus on the aspect of applications. Different from Larasati et al. [59], our paper gives a systematic, comprehensive survey on dataset distillation from a wide aspect of distillation frameworks, algorithms, applications, limitations, and promising directions.
@@ 3.1 Backpropagation Through Time Approach
$\nabla_{\mathcal{S}}\mathcal{L}(\mathcal{S})=\frac{\partial\mathcal{L}}{\partial\mathcal{S}}=\frac{\partial\mathcal{L}}{\partial\bm{\theta}^{(T)}}\left(\sum_{k=0}^{k=T}\frac{\partial\bm{\theta}^{(T)}}{\partial\bm{\theta}^{(k)}}\cdot\frac{\partial\bm{\theta}^{(k)}}{\partial\mathcal{S}}\right),\qquad\frac{\partial\bm{\theta}^{(T)}}{\partial\bm{\theta}^{(k)}}=\prod_{i=k+1}^{T}\frac{\partial\bm{\theta}^{(i)}}{\partial\bm{\theta}^{(i-1)}}.$ (9)
@@
Figure 5: (a) Meta-gradient backpropagation in BPTT [132]; (b) Meta-gradient backpropagation in kernel ridge regression; and (c) Meta-gradient backpropagation in gradient matching.
@@ 3.2 Kernel Ridge Regression Approach
where $\tau$ is a positive learned temperature scaling parameter, and $\hat{y}_{t}=\mathbf{K}_{X_{t}X_{s}}\left(\mathbf{K}_{X_{s}X_{s}}+\lambda I\right)^{-1}y_{s}$ is still calculated using the KRR formula.
@@ 4.1 Gradient Matching Approach
(a) Trajectory matching
(b) Distribution matching
Figure 6: (a) An illustration of trajectory matching [13]; (b) Compared to gradient matching approach, distribution matching approach can more comprehensively cover the data distribution in feature space.
@@ 4.3 Distribution Matching Approach
Therefore, the total loss in [113] for dataset distillation is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{f}}+\beta\mathcal{L}_{\text{d}}$, where $\beta$ is a positive factor for balancing the feature matching and discrimination loss.
@@ 5 Factorized Dataset Distillation
(a) Nonfactorized dataset distillation
(b) Factorized dataset distillation
Figure 8: Schematic diagrams of nonfactorized dataset distillation and factorized dataset distillation.
Code-based DD aims to learn a series of low-dimensional codes for generating highly informative images through a specific generator. In [127], the vectors that are put into the GAN generator are learned to producing informative images. Concretely, these investigators inverse real examples with a GAN generator and collect corresponding latent codes. Then, the latent vectors are further optimized with the distribution matching algorithm. By this, the optimized latent vectors can induce more informative synthetic examples with the pretrained GAN generator. In addition to use the GAN, Kim et al. [54] employed a deterministic multi-formation function $\texttt{multi-form}(\cdot)$ as the decoder to create synthetic data from fewer condensed data, i.e., the synthetic data are generated by $\mathcal{S}=\texttt{multi-form}(\mathcal{C})$. Then the condensed data $\mathcal{C}$ is optimized in an end-to-end fashion by the gradient matching approach.
@@
where the multiplication is element-wise operation. In addition, to enlarge the knowledge divergence encoded by different hallucinator networks for efficient compression, these researchers proposed an adversarial contrastive constraint that maximizes the divergence of features of different generated synthetic data. A similar code-decoder factorization method is also presented in Lee et al. [62], where they adopted an improved distribution matching to optimize the latent codes and decoders.
@@ 6 Performance Comparison
Table 1: Performance comparison of different dataset distillation methods on different datasets. Abbreviations of GM, TM, DM are for gradient matching, trajectory matching, and distribution matching, respectively. Whole denotes the test accuracy in terms of the whole target dataset.
@@ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$
$\mathbf{43.4}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
$50.7$
@@
$\mathbf{10.4}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{24.5}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ -
@@
$\mathbf{66.4}$ $\scriptscriptstyle\mathbf{\pm 0.8}$ $\mathbf{72.6}$ $\scriptscriptstyle\mathbf{\pm 0.7}$
@@
$\mathbf{85.0}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ $\mathbf{49.9}$ $\scriptscriptstyle\mathbf{\pm 0.2}$
@@
FRePo [132] KRR $\mathbf{93.0}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{98.6}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ $\mathbf{99.2}$ $\scriptscriptstyle\mathbf{\pm 0.0}$ $\mathbf{75.6}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
@@
$\mathbf{28.7}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ $\mathbf{42.5}$ $\scriptscriptstyle\mathbf{\pm 0.2}$
@@
RFAD [74] KRR $\mathbf{94.4}$ $\scriptscriptstyle\mathbf{\pm 1.5}$ $\mathbf{98.5}$ $\scriptscriptstyle\mathbf{\pm 0.1}$
@@
$\mathbf{78.6}$ $\scriptscriptstyle\mathbf{\pm 1.3}$ $\mathbf{87.0}$ $\scriptscriptstyle\mathbf{\pm 0.5}$ $\mathbf{88.8}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{52.2}$ $\scriptscriptstyle\mathbf{\pm 2.2}$
@@
IDC [54] GM - - - - - -
@@
$\mathbf{47.0}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ - - -
@@
Table 2: Performance comparison of different dataset distillation methods on ImageNet-1K.
@@
TESLA [25] $\mathbf{7.7}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{10.5}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{17.8}$ $\scriptscriptstyle\mathbf{\pm 1.3}$ $\mathbf{27.9}$ $\scriptscriptstyle\mathbf{\pm 1.2}$
@@
Table 3: Cross-architecture performance w.r.t. ResNet-18 (RN-18) and ResNet-152 (RN-152) on CIFAR-10. The results of DC, DSA, MTT, DM, and KIP are derived from Cui et al. [24]. The results of FTD and CAFE are collected from Du et al. [29]; and the result of TESLA is derived from Cui et al. [25].
@@
ConvNet RN-18 RN-152 ConvNet RN-18 RN-152 ConvNet RN-18 RN-152
@@
$\mathbf{15.3}$ $\scriptscriptstyle\mathbf{\pm 0.4}$

### v3

@@ Abstract
Deep learning technology has developed unprecedentedly in the last decade and has become the primary choice in many application domains. This progress is mainly attributed to a systematic collaboration in which rapidly growing computing resources encourage advanced algorithms to deal with massive data. However, it has gradually become challenging to handle the unlimited growth of data with limited computing power. To this end, diverse approaches are proposed to improve data processing efficiency. Dataset distillation, a dataset reduction method, addresses this problem by synthesizing a small typical dataset from substantial data and has attracted much attention from the deep learning community. Existing dataset distillation methods can be taxonomized into meta-learning and data matching frameworks according to whether they explicitly mimic the performance of target data. Although dataset distillation has shown surprising performance in compressing datasets, there are still several limitations such as distilling high-resolution data or data with complex label spaces. This paper provides a holistic understanding of dataset distillation from multiple aspects, including distillation frameworks and algorithms, factorized dataset distillation, performance comparison, and applications. Finally, we discuss challenges and promising directions to further promote future studies on dataset distillation.
Index Terms:
Efficient Deep Learning, Neural Network, Data Compression, Dataset Distillation
@@ 1 Introduction
During the past few decades, deep learning has achieved remarkable success due to powerful computing resources [1, 2, 3, 4, 5], which allow deep neural networks to directly handle giant datasets and bypass complicated manual feature extraction [6, 7]. For example, the powerful large language model GPT-3 contains $175$ billion parameters and is trained on $45$ terabytes of text data with thousands of graphics processing units (GPUs) [8]. However, massive data are generated every day [9], which pose a significant threat to training efficiency and data storage, and deep learning gradually reaches a bottleneck due to the mismatch between the volume of data and computing resources [10]. The emergence of dataset distillation (DD) is precisely to address this problem of large data volume. Dataset distillation, which was first proposed by [11] as a data compression technique that synthesizes a few high-informative data points to summarize numerous real data. Then, downstream models trained on the small synthetic dataset can achieve comparable generalization performance to those trained on the real data. A general illustration for dataset distillation is presented in Figure 1.
Before the appearance of DD, coreset selection [12, 13, 14] plays a pivotal role in dataset reduction by selecting a few prototype examples from the original training dataset as the coreset, and then the model is solely trained on the small coreset to save training costs while avoiding large performance drops. However, elements in the coreset are unmodified and constrained by the original data, which considerably restricts the coreset’s expressiveness, especially when the coreset budget is limited. Different from coreset selection, dataset distillation removes the restriction of uneditable elements and carefully modifies a small number of examples to preserve more information, as shown in Figure 3 for synthetic examples. By distilling the knowledge of the large original dataset into a small synthetic set, models trained on the distilled dataset can acquire a better generalization performance compared to the uneditable coreset.
@@
Due to the property of extremely high dimensions in the deep learning regime, the data information is hardly disentangled to specific concepts, and thus distilling numerous high-dimensional data into a few points is not a trivial task. Based on the objectives applied to mimic target data, dataset distillation methods can be grouped into meta-learning framework (Sec. 3) and data matching framework (Sec. 4), and these techniques in each framework can be further classified in a much more detailed manner. In the meta-learning framework, the distilled data are considered hyperparameters and optimized in a nested loop fashion according to the distilled-data-trained model’s risk w.r.t. the target data [36, 11]. The data matching framework updates distilled data by imitating the influence of target data on model training from parameter or feature space [24, 27, 31]. Figure 2 presents these different categories of DD algorithms in a tree diagram.
Apart from directly considering the synthetic examples as optimization objectives, in some studies, a proxy model is designed, which consists of latent codes and decoders to generate highly informative examples and to resort to learning the latent codes and decoders (Sec. 5). For example, Such et al. [17] employed a network to generate highly informative data from noise and optimized the network with the meta-learning framework. Zhao and Bilen [33] optimized the vectors and put them into the generator of a well-trained generative adversarial network (GAN) to produce the synthetic examples. Moreover, Kim et al. [26], Deng and Russakovsky [18], Liu et al. [30], Lee et al. [35] learned a couple of latent codes and decoders, and then synthetic data were generated according to the different combinations of latent codes and decodes. With this factorization of synthetic data, the compression ratio of DD can be further decreased, and the performance can also be improved due to the intraclass information extracted by latent codes.
In this paper, we present a comprehensive survey on dataset distillation, and the main objectives of this survey are as follows: (1) present a clear and systematic overview of dataset distillation; (2) review the recent progress and state-of-the-art algorithms and discuss various applications in dataset distillation; (3) give a comprehensive performance comparison w.r.t. different dataset distillation algorithms; and (4) provide a detailed discussion on the limitation and promising directions of dataset distillation to benefit future studies. We note that there is a concurrent survey on dataset distillation [37]. The most notable difference between [37] and our survey is the taxonomy of DD. It plainly classifies DD algorithms into four different categories of meta-model matching, gradient matching, trajectory matching, and distribution matching, while our hierarchical taxonomy first introduces meta-learning and data matching frameworks, which are then followed by a more fine-grained division. Therefore, our taxonomy provides a more systematic framework to track and summarize the development of DD algorithms and better sheds light on future DD studies.
@@ 2.1 Formalizing Dataset Distillation
Figure 4: (a) Meta-gradient backpropagation in BPTT [21]; (b) Meta-gradient backpropagation in kernel ridge regression; and (c) Meta-gradient backpropagation in gradient matching. The gradient $g_{\mathcal{S}}^{(T)}=\nabla_{\bm{\theta}^{(T)}}\mathcal{R}_{\mathcal{S}}(\bm{\theta}^{(T)})$.
@@ 3.1 Backpropagation Through Time Approach
$\nabla_{\mathcal{S}}\mathcal{L}(\mathcal{S})=\frac{\partial\mathcal{L}}{\partial\mathcal{S}}=\frac{\partial\mathcal{L}}{\partial\bm{\theta}^{(T)}}\left(\sum_{k=0}^{k=T}\frac{\partial\bm{\theta}^{(T)}}{\partial\bm{\theta}^{(k)}}\cdot\frac{\partial\bm{\theta}^{(k)}}{\partial\mathcal{S}}\right),$ (9)
and
$\frac{\partial\bm{\theta}^{(T)}}{\partial\bm{\theta}^{(k)}}=\prod_{i=k+1}^{T}\frac{\partial\bm{\theta}^{(i)}}{\partial\bm{\theta}^{(i-1)}}.$ (10)
@@ 3.2 Kernel Ridge Regression Approach
where $\tau$ is a positive learned temperature scaling parameter, and the prediction $\hat{y}_{t}=\mathbf{K}_{X_{t}X_{s}}\left(\mathbf{K}_{X_{s}X_{s}}+\lambda I\right)^{-1}y_{s}$ is still calculated using the KRR formula.
@@ 4 Data Matching Framework
(a) Trajectory matching
(b) Distribution matching
Figure 5: (a) An illustration of trajectory matching [27]; (b) T-SNE [46] visualization on the first class of CIFAR-10. Compared to gradient matching (Grad Match, marked with purple triangles), distribution matching (Dist Match, marked with red stars) can more comprehensively cover the real data distribution in feature space.
@@ 4.3 Distribution Matching Approach
Therefore, the total loss in [32] for dataset distillation is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{f}}+\beta\mathcal{L}_{\text{d}}$, where $\beta$ is a positive factor for balancing the feature matching and discrimination loss. Similar to the discrimination loss, Zhao et al. [34] add a classification loss $\mathcal{L}_{CE}$ as regularization to mitigate less classified synthetic data that caused by the first-order moment mean matching:
$\mathcal{L}_{CE}(\mathcal{S})=\sum_{i=1}^{|\mathcal{S}|}CE(f_{\bm{\theta}}(s_{i}),y_{i}),$ (36)
where the network parameter $\bm{\theta}$ is varied in different loops.
@@ 4.4 Discussion
Figure 7: Schematic diagrams of nonfactorized dataset distillation and factorized dataset distillation. Unlike directly updating synthetic (Syn) images in nonfactorized $D$, factorized DD optimizes the codes and decoders, which are then combined to generate synthetic data.
@@ 5 Factorized Dataset Distillation
Code-based DD aims to learn a series of low-dimensional codes for generating highly informative images through a specific generator. In [33], the vectors that are put into the GAN generator are learned to producing informative images. Concretely, these investigators inverse real examples with a GAN generator and collect corresponding latent codes. Then, the latent vectors are further optimized with the distribution matching algorithm. By this, the optimized latent vectors can induce more informative synthetic examples with the pretrained GAN generator. In addition to use the GAN, Kim et al. [26] employed a deterministic multi-formation function $\texttt{multi-form}(\cdot)$ as the decoder to create synthetic data from fewer condensed data, i.e., the synthetic data are generated by $\mathcal{S}=\texttt{multi-form}(\mathcal{C})$. Then the condensed data $\mathcal{C}$ is optimized in an end-to-end fashion by the gradient matching approach. Besides, the Partitioning and Expansion augmentation strategy adopted in [34] can also be regarded as a non-parametric decoder to generate more synthetic data based on the distilled codes. Concretely, while the distilled codes possess the same dimension as real images, Zhao et al. [34] firstly partition each code into several patches and then expand the small patches into the standard dimension during both dataset distillation and network training periods. Recently, Cazenavette et al. [53] argued that this low cross-architecture transferability comes from direct optimization on the pixel space, which makes the distilled data overfit to specific architectures used in the DD process. With this argument, they regularized the distillation by optimizing the intermediate latent codes that are fed into pretrained generative models for synthetic data generation, which largely increases the cross-architecture performance due to the more realistic synthetic data.
@@
where the multiplication is element-wise operation. In addition, to enlarge the knowledge divergence encoded by different hallucinator networks for efficient compression, these researchers proposed an adversarial contrastive constraint that maximizes the divergence of features of different generated synthetic data. A similar code-decoder factorization method is also presented in Lee et al. [35], where they adopted an improved distribution matching to optimize the latent codes and decoders. Recently, Zhang et al. [54] employ the powerful generative model as the decoder and optimize it and corresponding latent codes. Specifically, each low-dimensional latent code can generate synthetic images with different classes by feeding the generative model with corresponding conditional class embedding. By this, the number of learnable parameters is not affected by both the real image resolution and the number of classes, which is promising for tackling the large-scale distillation problem.
@@ 6 Performance Comparison
TABLE I: Performance comparison of different dataset distillation methods on different datasets. Abbreviations of GM, TM, DM are for gradient matching, trajectory matching, and distribution matching, respectively. Whole denotes the test accuracy in terms of the whole target dataset. The factorized DD methods are marked by .
@@ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$
$\mathbf{43.4}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{50.7}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
$10.4$
@@
$24.5$
$\scriptscriptstyle\pm 0.2$
@@
$66.4$
$\scriptscriptstyle\pm 0.8$
$72.6$
$\scriptscriptstyle\pm 0.7$
@@
$\mathbf{85.0}$ $\scriptscriptstyle\mathbf{\pm 0.1}$
$49.9$
$\scriptscriptstyle\pm 0.2$
@@
FRePo [21] KRR
$93.0$
$\scriptscriptstyle\pm 0.4$
$\mathbf{98.6}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ $\mathbf{99.2}$ $\scriptscriptstyle\mathbf{\pm 0.0}$
$75.6$
$\scriptscriptstyle\pm 0.3$
@@
$\mathbf{28.7}$ $\scriptscriptstyle\mathbf{\pm 0.1}$
$42.5$
$\scriptscriptstyle\pm 0.2$
@@
RFAD [22] KRR $\mathbf{94.4}$ $\scriptscriptstyle\mathbf{\pm 1.5}$
$98.5$
$\scriptscriptstyle\pm 0.1$
@@
$\mathbf{78.6}$ $\scriptscriptstyle\mathbf{\pm 1.3}$
$87.0$
$\scriptscriptstyle\pm 0.5$
$88.8$
$\scriptscriptstyle\pm 0.4$
$\mathbf{52.2}$ $\scriptscriptstyle\mathbf{\pm 2.2}$
@@
RCIG [23] KRR $\mathbf{94.7}$ $\scriptscriptstyle\mathbf{\pm 0.5}$ $\mathbf{98.9}$ $\scriptscriptstyle\mathbf{\pm 0.0}$ $\mathbf{99.2}$ $\scriptscriptstyle\mathbf{\pm 0.0}$ $\mathbf{79.8}$ $\scriptscriptstyle\mathbf{\pm 1.1}$ $\mathbf{88.5}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{90.2}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ - - - $\mathbf{53.9}$ $\scriptscriptstyle\mathbf{\pm 1.0}$ $\mathbf{69.1}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{73.5}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{39.3}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{44.1}$ $\scriptscriptstyle\mathbf{\pm 0.4}$
$46.7$
$\scriptscriptstyle\pm 0.3$
$\mathbf{25.6}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{29.4}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ -
IDC [26] GM
$94.2$
$\scriptscriptstyle\pm 0.2$
$98.4$
$\scriptscriptstyle\pm 0.1$
$99.1$
$\scriptscriptstyle\pm 0.1$
$81.0$
$\scriptscriptstyle\pm 0.2$
$86.0$
$\scriptscriptstyle\pm 0.3$
$86.2$
$\scriptscriptstyle\pm 0.2$
@@
DREAM [55] GM
$95.7$
$\scriptscriptstyle\pm 0.3$
$98.6$
$\scriptscriptstyle\pm 0.1$
$99.2$
$\scriptscriptstyle\pm 0.1$
$81.3$
$\scriptscriptstyle\pm 0.2$
$86.4$
$\scriptscriptstyle\pm 0.3$
$86.8$
$\scriptscriptstyle\pm 0.3$
$69.8$
$\scriptscriptstyle\pm 0.8$
$87.9$
$\scriptscriptstyle\pm 0.4$
$90.5$
$\scriptscriptstyle\pm 0.1$
$51.1$
$\scriptscriptstyle\pm 0.3$
$69.4$
$\scriptscriptstyle\pm 0.4$
$74.8$
$\scriptscriptstyle\pm 0.1$
$29.5$
$\scriptscriptstyle\pm 0.3$
$46.8$
$\scriptscriptstyle\pm 0.7$
$\mathbf{52.6}$ $\scriptscriptstyle\mathbf{\pm 0.4}$
$10.0$
$\scriptscriptstyle\pm 0.4$
$23.9$
$\scriptscriptstyle\pm 0.4$
$\mathbf{29.5}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
@@
$47.0$
$\scriptscriptstyle\pm 0.2$
IDM [34] DM - - - - - - - - -
$45.6$
$\scriptscriptstyle\pm 0.7$
$58.6$
$\scriptscriptstyle\pm 0.1$
$67.5$
$\scriptscriptstyle\pm 0.1$
$20.1$
$\scriptscriptstyle\pm 0.3$
$45.1$
$\scriptscriptstyle\pm 0.1$
$50.0$
$\scriptscriptstyle\pm 0.2$
$10.1$
$\scriptscriptstyle\pm 0.2$
$21.9$
$\scriptscriptstyle\pm 0.2$
$27.7$
$\scriptscriptstyle\pm 0.3$
@@
TABLE II: Performance comparison of different dataset distillation methods on ImageNet-1K.
@@
TESLA [29]
$7.7$
$\scriptscriptstyle\pm 0.2$
$10.5$
$\scriptscriptstyle\pm 0.3$
$\mathbf{17.8}$ $\scriptscriptstyle\mathbf{\pm 1.3}$ $\mathbf{27.9}$ $\scriptscriptstyle\mathbf{\pm 1.2}$
RCIG [23] $\mathbf{15.6}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{16.6}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ - -
@@
TABLE III: Cross-architecture performance w.r.t. ResNet (RN), DenseNet (DN), and ViT on CIFAR-10. The factorized DD methods are marked by .
@@
ConvNet RN-10 RN-18 RN-152 DN-121 ViT ConvNet RN-10 RN-18 RN-152 DN-121 ViT ConvNet RN-10 RN-18 RN-152 DN-121 ViT
@@
$\mathbf{15.3}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ - $\mathbf{24.68}$ $\scriptscriptstyle\mathbf{\pm 0.7}$